Aspire's Library

A Place for Latest Exam wise Questions, Videos, Previous Year Papers,
Study Stuff for MCA Examinations - NIMCET

Previous Year Question (PYQs)



Consider a domain consisting of three Boolean variables Toothache, Cavity, and Catch. The full joint distribution is a $2\times2\times2$ table as shown below.
toothache$\neg$toothache
catch$\neg$catchcatch$\neg$catch
cavity$0.108$$0.012$$0.072$$0.008$
$\neg$cavity$0.016$$0.064$$0.144$$0.576$
The probability of a cavity, given evidence of a toothache, $P(cavity\mid toothache)$ is ________.





Solution

We need:

$P(cavity\mid toothache)=\frac{P(cavity\cap toothache)}{P(toothache)}$

From table,

$P(cavity\cap toothache)=0.108+0.012=0.120$

Also,

$P(toothache)=0.108+0.012+0.016+0.064=0.200$

So,

$P(cavity\mid toothache)=\frac{0.120}{0.200}=0.600$


Online Test Series,
Information About Examination,
Syllabus, Notification
and More.

Click Here to
View More


Online Test Series,
Information About Examination,
Syllabus, Notification
and More.

Click Here to
View More

Ask Your Question or Put Your Review.

loading...